Planning #3 (similar to): Future Value | |- Question Help Future Value. Luis wants to know how much he will have available to spend on his trip to Belize in three years if he deposits $1,900 today at an interest rate of 11%. If he deposits $1,900 today at an annual interest rate of 11 percent, the amount Luis will have available to spend on his trip to Belize in three years is s. (Round to the nearest cent.) (Use your financial calculator or you may use the Financial Tables in Appendix C in computing your answer.)
Answer #1

By compound interest formula:

Amount = P*(1+r)n

A = 1900 * 1.113 = $2598.5

Please do rate me and mention doubts, if any, in the comments section.